OpenStudy (anonymous):

I need help with my homework graph the liner equations y-3=2(x-4)

OpenStudy (britanyashton):

okay. I'll help you through the equation. :) First, do this for me. Distribute the parenthesis on the right side soo expand it :) I'll wait, and then tell me what u get and ill check it.

OpenStudy (britanyashton):

After that, we'll add* the three to each side of the formula

OpenStudy (britanyashton):

We need to do this first, so that we can graph it easily.

OpenStudy (britanyashton):

Lemme know if you're here right now.

OpenStudy (anonymous):

ok sorry i was on another problem lol I got y-3=2x-8

OpenStudy (britanyashton):

Ok :) Perfect!! now add three to each side. You know why we do this right? Basic algebra. We want to make this equation in standard form. y=mx+b. Do that now!

OpenStudy (anonymous):

ok I am done

OpenStudy (britanyashton):

Now is it mainly the graphing part that confuses u? Do you still need me to explain that part?

OpenStudy (anonymous):

yes I was not sure how to get the second set

OpenStudy (britanyashton):

Ok. I got your back here. This confused me like crazy in Algebra too! y=mx+b m is the slope soo rise over run. Which means rise/run. Rise means up or down. So if you go up on a plane, your going a positive number, like 1+. If you go down, you are going down a negative amount of times, so it will be a negative number like -1, ect. We figure this out by picking ANY to points on the graph we make, and counting the rise/run. Since we're using a fraction, we can simplify it like any other one, and that fraction will also be true. :) B is the Y intercept. Which is the point where the line will intercept the y axis (the vertical line). Lemme show you, using the answer we just got. y-3=2x-8 +3 +3 y=2x-5 2 is the slope. Which is actually 2/1, which is two rise, and 1 run. So 2 up, one right. that makes sense right? But before we do that first we graph the y intercept! Our y intercept is -5. So graph that point.